OpenStudy (anonymous):

need someone to walk me through this question, please:

OpenStudy (anonymous):

Use the appropriate rule or combination of rules to find the derivative of the function defined below. \[y=2(\sin(6x)+\cos(x^3))^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

not even sure how to start this one. ;( looks like it would the the chain rule of course.

hartnn (hartnn):

first what is derivative of x^{1/2} = ?

OpenStudy (anonymous):

1/2x^-1/2

hartnn (hartnn):

we always go from outer function to inner function

hartnn (hartnn):

\(y'=2(1/2)(\sin(6x)+\cos(x^3))^\frac{ -1 }{ 2 }d/dx(\sin(6x)+\cos(x^3))\)

OpenStudy (anonymous):

options for the right answer are: \[a)2*\frac{ \cos(6x)-\sin(x^3) }{ (\sin(6x)+cox(x^3))^1/2 }\] \[b)\frac{ \cos(6x) +\sin(x^3)}{ (\sin(6x)+\cos(x^3))^\frac{ 1 }{ 2 } }\] \[c)\frac{ 12 }{ 6x^\frac{ 1 }{ 2 } }*\cos(6x)-\frac{ 6 }{ \cos(x^3)^\frac{ 1 }{ 2 } }*\sin(x^3)*x^2\] \[d)\frac{ 1 }{ (\sin(6x)^\frac{ 1 }{ 2 }+\cos(x^3))^\frac{ 1 }{ 2 } }\] \[e)\frac{ 6*\cos(6x)-3x^2*\sin(x^3) }{ (\sin(6x)+\cos(x^3))^\frac{ 1 }{ 2 } }\] f) same as e only 2 outside of the fraction. \[g) \frac{ 2 }{ \sin(6x)+\cos(x^3))^1/2 }\] \[h) \frac{ 6 }{ \sin(6x)^\frac{ 1 }{ 2 } }*\cos6x-\frac{ 3 }{ \cos(x^3)^\frac{ 1 }{ 2 } }*\sin(x^3)*x^2\]

hartnn (hartnn):

u got my last comment ? how i got that ?

OpenStudy (anonymous):

and i had that so far.

OpenStudy (anonymous):

yesi got that.

hartnn (hartnn):

so what is d/dx(sin(6x)+cos(x^3)) ?

OpenStudy (anonymous):

do i use the product rule for sin*6x and cos*x^3?

hartnn (hartnn):

there is no product, so NO u can treat them individually also what is d/dx (sin 6x) = ?

OpenStudy (anonymous):

cos6x

OpenStudy (anonymous):

-sinx^3

hartnn (hartnn):

only ?

hartnn (hartnn):

i mean only cos 6x ?

OpenStudy (anonymous):

times 6?

hartnn (hartnn):

u need chain rule here too, right ?

hartnn (hartnn):

yes! times 6, because u'll have d/dx(6x)

OpenStudy (anonymous):

cox6x*6

OpenStudy (anonymous):

cos6x*6

OpenStudy (anonymous):

????

hartnn (hartnn):

yup, what about 2nd term ? d/dx (cos (x^3)) = ?

OpenStudy (anonymous):

-sinx^3(3x^2)

hartnn (hartnn):

absolutely correct! u got it :)

hartnn (hartnn):

so which option is it ?

OpenStudy (anonymous):

e??

hartnn (hartnn):

put everything together, and u'll get it.

hartnn (hartnn):

yes, correct :)

OpenStudy (anonymous):

awesome!

hartnn (hartnn):

i hope chain rule is clearer to u now.....